OpenStudy (anonymous):

Solve 2 - 3x = x - 10

OpenStudy (whpalmer4):

Is that \[2-3x=x-10\] or \[2-3^x=x-10\]

OpenStudy (anonymous):

x=3

OpenStudy (anonymous):

if it is 2-3x=x-10

OpenStudy (anonymous):

a) x=-3 b) x=-2 c) x=2 d) x=3 those are my answers

OpenStudy (anonymous):

its d

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

because 2-3(3)=3-10 2-9=3-10 7=7

OpenStudy (anonymous):

no problem